Systems Staff Meeting Notes 02/25/02

Systems Staff Meeting Notes
February 25, 2002

Present: Rich Bennett, Will Chaney, Bill Covey, Suzy Covey, Debra Harris, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. Bobbie has a new grandson: Donaven Patrick – 7lbs. - 19 inches long. Has been quiet so far.
  2. Rich went to Silver Springs over the weekend and took advantage of the free admission for state employees. Saw Jim Nabors in concert, was a good show.
  3. Power outage this Sunday – we’ll need to turn off and unplug all power where possible – just like our hurricane preparedness drill. Battery UPS will remain plugged in. East north and south towers will be out. All of the Libraries will be closed that day. Rich will put a caution up on the Web site. Will notify the liaisons, explaining that the power surge when the power resumes is the reason everything must be unplugged.
Review Minutes of 2/11/02 and 2/18/02
  1. 2/11/02 Minutes – Review #5 – John Van Hook’s machine - HD has arrived, has been installed and is working.
  2. 2/18/02 Minutes – Equipment RIR #1 – purchasing wish list for Systems – need to get requests to Bill ASAP.
  3. 2/18/02 Minutes – Liaisons #7 – Default language problem resolved – just need to reboot.
  4. 2/18/02 Minutes – Networking #8 – Xerox meeting never happened (again). It was cancelled; Bill was not informed. Meeting has been rescheduled for tomorrow. Tech support and other issues are to be resolved.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. The 30 Dells arrived – only three have been allocated so far. Unfortunately, they require a different type of SDRAM then what we ordered. They take DDR, not PC133. We’ll have to buy different RAM this week.
  2. SAN order to go out this week – upgrade for NT2R. Second unit to go to MSL mailroom area rather than L206. Will, John Ashcraft and Carol Drum will meet today to firm up the space, then Facilities will need to work on the power, air conditioning and fire suppression. The fire suppression could be a big issue – Bill suggests a halon handheld unit.
  3. The Epson scanners have been received. They require a driver download to work with multiple users on W2K. The HP scanners need 2906 SCSI cards.
Support Issues and Liaisons
  1. Kiosks will run 3270 as the shell, but it wants the user to connect to a profile. Winston is working on an applet to launch the command line to open the profile. Default has no setting. Also want exit to restart the machine.
  2. FP2K disk – Pkg 6702 – is missing. Debra left it out on the work desk for Emily to pick up on Friday & now it is gone. Will check with Emily, Tom Minton & John Ashcraft.
Networking
  1. Individual mailbox backups are still working – with the one exception.
  2. Network browse function – the server on the 239 subnet did take over, but not all workstations are showing up. Need to make sure the local host file is correct. Each master browser needs to know who the PDC is.
  3. W2K server conversion – must have DNS.
  4. Centralized software installation – Bill is looking into it.
  5. Geoffrey has received more requests from super liaisons to make modifications to the public user profiles – they are currently locked down. If they can be stably converted we can give some super-liaisons the rights to modify, but they’ll need to understand that they will be the ones to fix any problems. Will need to make new user groups for each one. We’ll give it a try and see how it goes.
Software Projects Status
  1. DLC – Sanborn is the current project. Ying is writing the code for the data input. There is a set of maps that are confusing – the sheets aren’t on a grid pattern. Winston says it’s under control and does make sense - you just have to see it to understand it. Erich is trying to generalize it to accept searches from other applications – a Geographical Core. That has not been implemented yet.
  2. Hilton is finished with FilmLog. He is looking at PHP programs for ARIEL solutions. Winston needs to get with Suzy to put it on the Web server. Rich notes that Bill Koontz has the same database, he may let us use it and we may not need to worry about this.
  3. Cynthia is working on the ILL Content site. Will demo it this week.
  4. Winston tried to put the new version of SQL server on a test machine - NT14 – but it refused to load. Will put W2K advanced server on another box to test.
DP Coordinator
  1. Spring Break is next week: March 4-8. Students will start dropping off this Friday.
  2. Facilities picked up our old water cooler.


Back to the Systems Homepage

Last updated March 5, 2002
by Debra Harris
debharr@mail.uflib.ufl.edu